Sam dug a pit in his garden in a shape of a rectangular prism with 9 meters length and 5 meters width if the volume of soil dug out is 81 meters cubed what is the depth of the pit. In a form of a decimal
Answer by ankor@dixie-net.com(22740)

Sam dug a pit in his garden in a shape of a rectangular prism with 9 meters length and 5 meters width if the volume of soil dug out is 81 meters cubed what is the depth of the pit. In a form of a decimal.
:
Volume = length * width * depth
L * w * d = 81
L=9; w=5, V=81, find d
9 * 5 * d = = 81
45d = 81
d = 81/45
d = 1.8 meters is the depth